OpenStudy (anonymous):

Solve And Show Steps 8/X^2-16 Subtracted By 1/X-4 = 1

OpenStudy (anonymous):

\[[8\div (x^2-16)] - [1\div (x-4)] = 1\] Is that what you're trying to solve?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Ok, well first you have to get a common denominator so you can add fractions. Since x^2-16 is the same as (x+4)(x-4), the common denominator will be just that (as (x-4) is the denominator of the second fraction). So you'll want to multiply the second fraction by (x+4) so they'll both have that common denominator. This will give you\[8-(x+4)\div (x+4)(x-4)=1\]

OpenStudy (anonymous):

Now I'm a little stuck so I have to think about this...oh! Well once you simplify that you'll have x+4/(x+4)(x-4)-1 so the (x+4)'s cancel out.

OpenStudy (anonymous):

So 1/(x-4)=1, multiply both sides by (x-4) 1=x-4 x=5